OpenStudy (anonymous):

Solve the equation below for x. Round your answer to the nearest thousandth: 10=3^(x+5)-2 A. -2.738 B. -4.558 C. 2.262 D. 5.442

OpenStudy (anonymous):

condense the equation into two terms, 12 = e^(x+5). Ideally, we want to take the logarithm of each side on base 3. This gives Log3(12) = x+5. Using the change of base theorum to make it calculator appropriate, log10(12)/log10(3)=x+5. This gives 2.262 = x+5. subtract 5 from both sides gives -2.738

OpenStudy (anonymous):

Thank you!